THE SPECTRA AND DYNAMICS OF DIATOMIC MOLECULES

\underline{ROBERT W. FIELD}, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; and HÉLÈNE LEFEBVRE-BRION, Laboratoire de Photophysique Moléculaire, Université Paris–Sud, Orsay, France.

A revised and significantly expanded version of “Perturbations in the Spectra of Diatomic Molecules” (Academic, 1986) is now available from Elsevier. The chapter titles are: 1. Simple Spectra and Standard Experimental Techniques, 2. Basic Models, 3. Terms Neglected in the Born-Oppenheimer Approximation, 4. Methods of Deperturbation, 5. Interpretation of the Perturbation Matrix Elements, 6. Transition Intensities and Special Effects, 7. Photodissociation Dynamics, 8. Photoionization Dynamics, 9. Dynamics. Chapters 2-8 contain all of the material in “Perturbations… ” (corrected and updated), but major new topics have been added to Chapters 3, 6, 7, and 8. Examples include a more complete treatment of Hund’s cases, a discussion of electromagnetic field dressed adiabatic states, expansions of the chapters on predissociation and autoionization to include direct photofragmentation (dissociation and ionization) processes, and a discussion of coherent control. Chapters 1 and 9 are entirely new. The book concludes with a long section that extends the concepts, tools, and visualization techniques beyond diatomic molecules to larger molecules. This book is written for beginning graduate students, users of spectra and spectroscopic data, and for experts, both experimentalists and theorists.